Fermi National Laboratory


Accelerator Update

Friday November 5

Tevatron

Bo Low Beta Separators HLS Installation - 75% complete

D0 Low Beta Separators HLS Installation - 60% complete

IPM installation - 90% complete. The Crystal Collimator has been installed

D4 Spool Replacement - 80%

D17 Separator Installation - 99% complete

Shimming 412 Dipoles - 100% complete

Unrolling Magnet - 100% complete

Linac

The Linac is up and running.

Booster

Booster hit 4.2E12 and is running at 80% efficiency. NTF is coming up, but no patients have been scheduled.

Antiproton Source

The last section of the Debuncher Injection Area Beampipe will be installed next week. Work on the quads is finished.

AP1/3 Line -Vacuum pumpdown complete. Final work in progress. Expect to close up the Prevault and Pretarget area next week. D:H926 test to follow.

AP2 Line - The beam stop has been removed and the opening measured. The shunt/power supply reconfiguration nearly complete.

SO Cord Remediation - The AP50 work is in progress and will be completed next week.

D/A Lines - Alignment numbers in hand. A:ISEP1 has been removed for replacement.

Main Injector/Recycler

Overall shut down progress - 95% complete

MI-30

Q307-309 Vacuum Modifications- 40% complete

MI-30 & 62

Correction Magnet Replacement - overall ~ 95% complete

MI-10

Cooling Tank Repairs - ~95% complete

MiniBooNE

Waiting for parts. The experiment should be ready for beam by November 30.

NuMI

Most tasks finished. The horn has been successfully tested.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 11/8/2004   email Fermilab